OpenStudy (anonymous):

Determie whether the improper integral diverges or converges? Evaluate the integral if it converges.

OpenStudy (anonymous):

What integral?

OpenStudy (lgbasallote):

hi guys. been a long time since we all gathered together

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty}e ^{-x }\cos xdx\]

OpenStudy (experimentx):

find the indefinite integral first. then apply the limit.

OpenStudy (anonymous):

thanks so much guys

OpenStudy (experimentx):